x = 0 is an ordinary point of a certain linear differential equation. After the assumed solution y = ∞ n = 0 cnxn is substituted into the DE, the following algebraic system is obtained by equating the coefficients of x0, x1, x2, and x3 to zero.

2c2 + 2c1 + c0 = 0

6c3 + 4c2 + c1 = 0

12c4 + 6c3 + c2 − 1 3 c1 = 0

20c5 + 8c4 + c3 − 2 3 c2 = 0

Bearing in mind that c0 and c1 are arbitrary, write down the first five terms of two power series solutions of the differential equation.

y1(x) = c0 1 − 1 2 x2 + 1 3 x3 − 1 4 x4 + 1 5 x5 + and y2(x) = c1 x − x2 + 1 2 x3 − 1 3 x4 + 1 4 x5 +

y1(x) = c0 1 − x2 + 1 2 x3 − 5 36 x4 − 1 360 x5 + and y2(x) = c1 x − 1 2 x2 + 1 3 x3 − 1 8 x4 + 1 60 x5 +

y1(x) = c0 1 − 1 2 x2 + 1 3 x3 − 1 8 x4 + 1 60 x5 + and y2(x) = c1 x − x2 + 1 2 x3 − 5 36 x4 − 1 360 x5 +

y1(x) = c0 1 − 1 2 x2 + 1 6 x3 − 1 24 x4 + 1 120 x5 + and y2(x) = c1 x − 1 2 x2 + 1 6 x3 − 1 24 x4 + 1 120 x5 +

y1(x) = c0 1 − x2 + 1 2 x3 − 1 3 x4 + 1 8 x5 + and y2(x) = c1 x − x2 + 1 2 x3 − 5 36 x4 − 1 360 x5 +
Answer #1

Please like the answer, it will encourage me.

Consider the equations obtained in solving power series solution of DE: 2C+2C, +Co = 0 6C, +4C, +C =0 12C, + 6C, C =0 2003 +8168 C 20. 2 Cz = - +-- -- 9.20 .6.20 2.20 3.20 3.20° C3 = 40,+ 1c - C 1c 45° 15° 40 30 30° t-Co 36097 360 30 k=0 Therefore, t